find the eation of te circle which touches the yqu-axis at the origin and passes through the points (a,b)

Grade:Select Grade

1 Answers

Vikas TU
6 years ago
equation of the circle which touches the y-axis at the origin is given by
 
  • (0-h)^2 + ( 0 – k )^2 = r^2
  • h^2 + k^2 = r^2
equation of the circle which passes through (a,b) is given by
 
  • (a-h)^2 + ( b – k )^2 = r^2
  • a^2 + h^2 – 2ah + b^2 + k^2 – 2bk = r^2
  • Hence
 
a^2 + h^2 – 2ah + b^2 + k^2 – 2bk = h^2 + k^2
 
a^2 – 2ah + b^2 – 2bk = 0. Where (h,k) is the center.